a train travels 63mph for 20 minutes. How far does it travel?
Answer by jim_thompson5910(35256)

20 min = 20/60 = 1/3 hr

d = rt

d = 63(1/3)

d = 63/3

d = 21

So it travels 21 miles
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------